OpenStudy (anonymous):

Describe how the graphs of y = |x| and y = |x| – 15 are related. A. The graphs have the same shape. The y-intercept of y = |x| is 0, and the x-intercept of the second graph is –15. B. The graphs have the same y-intercept. The second graph is steeper than y = |x|. C. The two graphs are the same. D. The graphs have the same shape. The y-intercept of y = |x| is 0, and the y-intercept of the second graph is –15.

OpenStudy (anonymous):



OpenStudy (anonymous):

The first one is y=mod(x) and the second one is y=mod(x)-15. Compute values of y for x=0,-1,1,-2,2 and so on. You'll get the answer :)